OpenStudy (anonymous):

Given the function f(x) = x2 and k = 2, which of the following represents a horizontal shift? f(x) + k kf(x) f(x + k) f(kx)

OpenStudy (usukidoll):

f(x) = x^2 is the original graph and k = 2 which is the change

OpenStudy (usukidoll):

we graph the original f(x) first and then apply these transformations...

OpenStudy (isaiah.feynman):

A horizontal shift would be in the form \[f(x \pm k)\]

OpenStudy (usukidoll):

-_- wow Isaiah, you didn't have to give out a direct answer.

OpenStudy (karatechopper):

Just by rule.... f(x)+k = vertical shift kf(x) = vertical stretch/shrink (depending) f(x+k) = phase shift, also known as a horizantal shift. f(kx) = Horizantal stretch/shrink Those are the rules to always use.

OpenStudy (karatechopper):

I don't know how else to explain just by rule....forgive me @UsukiDoll if you or anyone thought this was my way of giving out an answer. Just thought I would give them a key to help them identify what they are handling with other problems.

OpenStudy (usukidoll):

@karatechopper your answer was fine.